Given real numbers a and b; let a*b = a+b+ab: Evaluate 3 * (4 * 5):
a) 60 (b) 80 (c) 119 (d) 133 (e) None of the above

Answer by bamigbe20(3)


Given real numbers a and b; let a*b = a+b+ab: Evaluate 3 * (4 * 5)
solution:
we solve this one at a time. i.e, we first evaluate the one with bracket, which is (4 * 5):
then, 4 * 5 = 4+5+4(5)
= 9+20
= 29
so, 4 * 5 = 29.

we then evaluate 3 * (4 * 5) , where 4 * 5 = 29.
then, 3 * 29 = 3+29+3(29)
= 3+29+87
= 119.
Therefore 3 * (4 * 5) = 119....................Hence, The correct answer option is "C".